OpenStudy (anonymous):

If each exterior angle or a regular polygon measures 15°, how many sides does the polygon have?

Parth (parthkohli):

Interior angle = 180 - 15 = 165 degrees

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {180(n - 2) \over n} = 165 }\) Solve for \(n\).

Parth (parthkohli):

I suppose you already knew the formula (180(n - 2))/2

Parth (parthkohli):

(180(n - 2))/n*